OpenStudy (anonymous):

Need help please. Write -2.375 as a mixed number in simplest form.

OpenStudy (anonymous):

-2 375/1000 = -2375/1000 now just simplify

OpenStudy (anonymous):

or just -2 with 375/100 = 3/8

OpenStudy (anonymous):

so -2 3/8

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."